### v1

@@ Abstract
Our findings show that all three models struggle to accurately detect and warn users about vulnerabilities, achieving a detection rate of only 12.6% to 40% across our datasets. We also observe that the LLMs tend to identify certain types of vulnerabilities related to sensitive information exposure and improper input neutralization much more frequently than other types, such as those involving external control of file names or paths. Furthermore, when LLMs do issue security warnings, they often provide more information on the causes, exploits, and fixes of vulnerabilities compared to Stack Overflow responses. Finally, we provide an in-depth discussion on the implications of our findings and present a CLI-based prompting tool that can be used to generate significantly more secure LLM responses.
@@
journal: Noname
@@ 1 Introduction
Beyond generating vulnerable code, using LLMs can impact software security in more intricate and subtle ways. For instance, novice developers may unknowingly input insecure code (copied from Q&A forums) and ask LLMs to refactor and adapt it to their problem context. Similarly, during debugging, a developer might provide a block of code containing vulnerabilities, such as unsanitized user input in an SQL query, without being aware of its potential security implications. If LLMs fail to identify and address these vulnerabilities, developers may integrate flawed code into their projects, relying on the model without recognizing the potential security risks themselves. To better understand this phenomenon, let us consider the following example. In a Stack Overflow question, a developer asks for help to resolve an issue with writing to a file:
@@
Here, the developer is focused on removing the extra %0A characters, which could be resolved using the strip() function i.e., requests.get(a.strip(), timeout=(2, 5), verify=False)). However, they fail to notice a significant security risk posed by the verify=False parameter. Disabling SSL certificate verification is generally considered poor security practice and exposes the application to serious risks, including man-in-the-middle attacks, where an attacker could intercept or manipulate the data sent over HTTPS. When prompted with this question, GPT-4, Claude 3, and Llama 3 all correctly explained the issue with %0A and suggested using the strip() function to fix it. However, none of the LLMs mentioned the security implications of the verify=False parameter, leaving the developer unaware of the potential vulnerability. By failing to inform the user of this vulnerability, the LLM responses can indirectly reinforce the faulty implementation and, in many cases, further build upon it, perpetuating insecure coding practices.
Several studies have explored the potential risks associated with the use of LLMs and examined concerns regarding the generation of insecure code (Pearce et al., 2022; Siddiq and Santos, 2022; Khoury et al., 2023; Siddiq et al., 2024b), inaccuracies in vulnerability detection (Ullah et al., 2024; Akuthota et al., 2023; Purba et al., 2023; Zhou et al., 2024b), and potential misuse for offensive applications, ranging from hardware to user-level attacks (Happe and Cito, 2023; Falade, 2023). However, most studies have focused on the security of LLM-generated code, often neglecting the natural language generated by LLMs that plays a critical role in interactive learning and problem-solving (Eastman, 2023; Xia and Zhang, 2023; Hao et al., 2024). Additionally, unlike prior research that focuses on detecting vulnerabilities in LLM-generated code or evaluating the capabilities of LLMs when explicitly tasked to detect vulnerabilities, our work investigates the ability of LLMs to proactively identify vulnerabilities in user-supplied code. This reflects real-world use cases where developers that rely on LLMs for various tasks inevitably prompt LLMs with vulnerable code. Our study is the first to assess the security awareness of LLMs using the textual information provided by the LLMs alongside or independent of code. We assess not only LLMs’ ability to proactively detect vulnerabilities but also their effectiveness in communicating critical information, such as causes, exploits, and fixes, to enhance developer understanding and prevent the use of exploitable and insecure code.
@@
  • A benchmark for evaluating: (a) LLMs’ capabilities in proactively detecting vulnerabilities in real-world user-supplied code, and (b) issuing warnings with critical information, such as causes, exploits, and fixes, to enhance developer understanding and mitigate integration of insecure code in existing code-bases.
@@ 2 Motivational Study
Types of Information: To examine the types of information in the ChatGPT conversations, we used the Pan et al. (2021) taxonomy for for categorizing information in developer chats. Developer chats (e.g., Slack, Discord) closely align with developer-LLM conversations since both support dynamic, rapid, and iterative information exchange (Chatterjee et al., 2019). Consequently, we believe this taxonomy is well-suited for our study. The information categories in the taxonomy are listed in Table
1.
@@
Given our observations, this empirical exploration of real-world developer conversations with ChatGPT highlights the LLM’s potential to proactively inform developers about security vulnerabilities and provide context-specific solutions for writing secure programs. Although infrequent, such proactive contributions promote security awareness, helping developers avoid insecure practices and prevent them from building upon vulnerable code. These findings highlighted the importance of systematically studying this behavior. Consequently, we designed our main study to evaluate the extent and consistency of this behavior across three popular LLMs, under controlled scenarios.
@@ 3.1 Dataset Collection and Refactoring
Data Collection. We collect a dataset of SO questions with insecure code but without any explicit mention of security within the questions themselves. Figure 3 shows an example of such a question, where the JavaScript code uses the eval function to dynamically create a variable based on user input, introducing a significant risk of code injection or a cross-site scripting (XSS) attack. The developer posting the question seems unaware of the related security issue; however, one of the respondents points this out in a comment “You have a pretty nasty XSS vulnerability here”.
@@ 5.2 Implications for Tool Integration in Software Development
Figure 9: Left: The original user prompt featuring vulnerable code. Right: The LLM’s response after prompt modification. Both the prompt and response have been abbreviated for presentation clarity.
We use the same 50 questions used in Section 5.1 to experiment with this approach. Having analyzed each question’s extracted code snippets with CodeQL, we add CodeQL’s description of the detected vulnerabilities to the end of each question, using the phrase “Address security vulnerability: <Vulnerability Description #1> Address security vulnerability <Vulnerability Description #2>…”. After prompting the GPT-4 with these prompts, the LLM responses were again analyzed with respect to RQ1 and RQ2. Here, similar to the previous section, the authors manually carried out the analysis.
@@
Several opportunities exist for integrating LLMs with SE tools in the development workflows. For instance, static analysis tool like CodeQL can be used to preprocess code snippets to outline specific security vulnerabilities in the prompt. The identified vulnerability can then be added as meta data to the prompt, providing the LLM with clearer directions on what issues need addressing. Such integration could enhance the LLM’s ability to generate secure code and streamline vulnerability detection and mitigation.
@@ 6 Threats to Validity
External Validity: The results of our study may not generalize to all LLMs or programming questions. For instance, the inclusion of other LLMs, such as code-specific models that place a heavier emphasis on code and provide more concise natural language output, is outside the scope of our study and can be explored further in future work. However, to investigate the security awareness of cutting-edge LLMs, we included three of the most popular models: GPT-4, Llama 3, and Claude 3. Further, we collected our datasets from a SO dump spanning 9 years (March 2015-March 2024). We selected questions related to two popular programming languages, Python and JavaScript, to ensure our findings adaptable to a wide range of development scenarios and questions.

### v2

@@ Abstract
Our findings show that all three models struggle to accurately detect and warn users about vulnerabilities, achieving a detection rate of only 12.6% to 40% across our datasets. We also observe that the LLMs tend to identify certain types of vulnerabilities related to sensitive information exposure and improper input neutralization much more frequently than other types, such as those involving external control of file names or paths. Furthermore, when LLMs do issue security warnings, they often provide more information on the causes, exploits, and fixes of vulnerabilities compared to Stack Overflow responses. Finally, we provide an in-depth discussion on the implications of our findings, and demonstrated a CLI-based prompting tool that can be used to produce more secure LLM responses.
@@
journal: Empirical Software Engineering
@@ 1 Introduction
Beyond generating vulnerable code, using LLMs can impact software security in more intricate and subtle ways. For instance, novice developers may unknowingly input insecure code (copied from Q&A forums) and ask LLMs to refactor and adapt it to their problem context. Similarly, during debugging, a developer might provide a block of code containing vulnerabilities, such as unsanitized user input in an SQL query, without being aware of its potential security implications. If LLMs fail to identify and address these vulnerabilities, developers may integrate flawed code into their projects, relying on the model without recognizing the potential security risks themselves. To better understand this phenomenon, let us consider the following examples. In a SO question, a developer asks for help to resolve an issue with writing to a file:
@@
Here, the developer is focused on removing the extra %0A characters, which could be resolved using the strip() function i.e., requests.get(a.strip(), timeout=(2, 5), verify=False)). However, they fail to notice a significant security risk posed by the verify=False parameter. Disabling SSL certificate verification is generally considered poor security practice and exposes the application to serious risks, including man-in-the-middle attacks, where an attacker could intercept or manipulate the data sent over HTTPS. If prompted with this question, GPT-4, Claude 3, and Llama 3 all correctly explain the issue with %0A and suggested using the strip() function to fix it. However, none of the LLMs mentioned the security implications of the verify=False parameter, leaving the developer unaware of the potential vulnerability. By failing to inform the user of this vulnerability, the LLM responses can indirectly reinforce the faulty implementation and, in many cases, further build upon it, perpetuating insecure coding practices.
Several studies have explored the potential risks associated with the use of LLMs and examined concerns regarding the generation of insecure code (Pearce et al., 2022; Siddiq and Santos, 2022; Khoury et al., 2023; Siddiq et al., 2024b), inaccuracies in vulnerability detection (Ullah et al., 2024; Akuthota et al., 2023; Purba et al., 2023; Zhou et al., 2024b), and potential misuse for offensive applications, ranging from hardware to user-level attacks (Happe and Cito, 2023; Falade, 2023). However, most studies have focused on the security of LLM-generated code, often neglecting the natural language generated by LLMs that plays a critical role in interactive learning and problem-solving (Eastman, 2023; Xia and Zhang, 2023; Hao et al., 2024). Additionally, unlike prior research that focuses on detecting vulnerabilities in LLM-generated code or evaluating the capabilities of LLMs when explicitly tasked to detect vulnerabilities, our work investigates the ability of LLMs to proactively identify vulnerabilities in user-supplied code. This reflects real-world use cases where developers that rely on LLMs for various tasks inevitably prompt LLMs with vulnerable code. Our study is the first to assess the security awareness of LLMs using the textual information provided by the LLMs alongside or independent of code. We assess not only LLMs’ ability to proactively detect vulnerabilities but also their effectiveness in communicating critical information – such as causes, exploits, and fixes – to enhance developer understanding and prevent the use of exploitable and insecure code.
@@
  • A benchmark for evaluating: (a) LLMs’ capabilities in proactively detecting vulnerabilities in real-world user-supplied code, and (b) issuing warnings with critical information—such as causes, exploits, and fixes—to enhance developer understanding and mitigate integration of insecure code in existing code-bases.
@@ 2 Motivational Study
Types of Information: To examine the types of information in the ChatGPT conversations, we used the Pan et al. (2021) taxonomy for for categorizing information in developer chats. Developer chats (e.g., Slack, Discord) closely align with developer-LLM conversations since both support dynamic, rapid, and iterative information exchange (Chatterjee et al., 2019). Consequently, we believe this taxonomy is well-suited for our study. The information categories in the taxonomy are listed in Table 1.
@@
Given our observations, this empirical exploration of real-world developer conversations with ChatGPT highlights its potential to proactively inform developers about security vulnerabilities and provide context-specific solutions for writing secure programs. Although infrequent, such proactive contributions promote security awareness, helping developers avoid insecure practices and prevent them from building upon vulnerable code. These findings highlighted the importance of systematically studying this behavior. Consequently, we designed our main study to evaluate the extent and consistency of this behavior across three popular LLMs, under controlled scenarios.
@@ 3.1 Dataset Collection and Refactoring
Data Collection.
We collect a dataset of SO questions with insecure code but without any explicit mention of security within the questions themselves. Figure 3 shows an example of such a question, where the JavaScript code uses the eval function to dynamically create a variable based on user input, introducing a significant risk of code injection or a cross-site scripting (XSS) attack. The developer posting the question seems unaware of the related security issue; however, one of the respondents points this out in a comment “You have a pretty nasty XSS vulnerability here”.
@@ 5.2 Implications for Tool Integration in Software Development
Figure 9: Comparison of the LLM’s original functionality-focused response (left) and its subsequent security-focused response (right). On the right, we observe that explicit mention of vulnerability caused the model to overlook the main question.
We use the same 50 questions used in Section 5.1 to experiment with this approach. Having analyzed each question’s extracted code snippets with CodeQL, we add CodeQL’s description of the detected vulnerabilities to the end of each question, using the phrase “Address security vulnerability: <Vulnerability Description #1> Address security vulnerability: <Vulnerability Description #2>…”. After prompting the GPT-4 with these prompts, the LLM responses were again analyzed with respect to RQ1 and RQ2. Here, similar to the previous section, the authors manually carried out the analysis.
@@
We also examined whether modifying the prompts affected the overall functional correctness of the LLM’s answers. In 46 out of the 50 responses, GPT-4 adequately addressed the user’s primary concern while fixing and/or acknowledging the security vulnerabilities. However, in 4 instances, the LLM’s response only addressed the security issue and completely overlooked the user’s main query. Figure 9 illustrates one such scenario where the question seeks guidance on preventing multiple database entries for the same user. Note that the user’s code insecurely handles logging credentials, introducing a potential security risk. Without a security-specific prompt, the LLM addresses the core issue, ignoring the vulnerability (Figure 9 - left panel). After integrating the CodeQL findings into the prompt, the LLM only addresses the logging and credential handling vulnerabilities while neglecting the main question (Figure 9 - right panel). This example demonstrates how emphasizing security can sometimes overshadow the original problem and highlights the challenge of balancing these objectives in real-world usage.
Despite the challenges, several opportunities exist for integrating LLMs with SE tools in the development workflows. For instance, static analysis tool like CodeQL can be used to preprocess code snippets to outline specific security vulnerabilities in the prompt. The identified vulnerability can then be added as meta data to the prompt, providing the LLM with clearer directions on what issues need addressing. Such integration could enhance the LLM’s ability to generate secure code and streamline vulnerability detection and mitigation.
@@
Figure 10: Left: The original user prompt featuring vulnerable code. Right: The LLM’s response after prompt modification (integrating CodeQL feedback). Both the prompt and response have been abbreviated for presentation clarity.
@@ 6 Threats to Validity
External Validity: The results of our study may not generalize to all LLMs or programming questions. One key limitation of our study is the exclusion of code-specific LLMs, such as DeepSeek-Coder Team (2024) and Qwen-Coder Hui et al. (2024), which are designed specifically for code-related tasks. These models place a stronger emphasis on code comprehension and may yield different security-related responses compared to the conversational LLMs we evaluated. While our focus was on conversational LLMs widely used in developer interactions, we acknowledge that analyzing the security awareness of code-specific models is an important avenue for future research. In investigating the security awareness of conversational LLMs, however, we included three of the most popular models: GPT-4, Llama 3, and Claude 3. Further, we collected our datasets from a SO dump spanning 9 years (March 2015–March 2024). We selected questions related to two popular programming languages, Python and JavaScript, to ensure our findings remain adaptable to a wide range of development scenarios and questions.
@@ 8 Conclusions
Compliance with Ethical Standards
Funding: Not applicable to this study.
Ethical approval: Not applicable to this study.
Informed consent: Not applicable to this study.
Author Contributions: Not applicable to this study.
@@
Conflict of interest: The authors declare that they have no conflict of interest.